Find the equation of a line passing through the pair of points. Write the equation in slope intercept form. The points are (2,1) and (10,-31)
Answer by Alan3354(69443)

Find the equation of a line passing through the pair of points. Write the equation in slope intercept form. The points are (2,1) and (10,-31)
------------
The hard way:
Find the slope, m = diffy/diffx = (-31-1)/(10-2) = -32/8
m = -4
---------
Use y = mx+b and either point to find b
1 = -4*2 + b
b = 9
--> y = -4x + 9
-------------
Easier way:
---
|+x +y +1|
|+2 +1 +1| = 0
|10 -31 1|
---------
x(1+31) - y(2-10) + (-62-10) = 0
32x + 8y = 72
4x + y = 9
y = -4x + 9